I just need to check my answers:

17: You want to buy a house in 5 years and expect to need $35000 for a down payment. If you have $11000 to invest, how much interest do you have to earn (compounded annually) to reach your goal? (Enter your answers as a decimal rounded to 4 decimal places, not a percentage. For example, enter 0.0843 instead of 8.43%)

18: You decide that the loan in question 2 is too large for you. You know you can afford to make monthly payments of $325 for 60 months. How much will the bank lend you today on a 3% APR car loan in exchange for your promised monthly payments?
Answer #1

17). r = [FV / PV]1/n - 1

= [$35,000 / $11,000]1/5 - 1 = 1.2605 -1 = 0.2605

18). PVA = Annuity x [{1 - (1 + r)-n} / r]

= $325 x [{1 - (1 + 0.03 / 12)-60} / (0.03 / 12)]

= $325 x [0.1391 / 0.0025] = $325 x 55.6524 = $18,087.02